The towing capacities (in pounds) of all the pickup trucks at a dealership have a
bell-shaped distribution, with a mean of 11,830 pounds and a standard deviation
of 2370 pounds. Use the corresponding z-score to determine
whether the towing capacity is unusual. Explain your reasoning.
46. 5500 pounds
Answer #1

Answer

we have mean = 11830 and standard deviation = 2370

we have to find the z score corresponding to x value of 5500

We know the formula for the z score is given as

z = (ar{x}-mu)/sigma

where we have -= 5500, μ-1 1 830 σ-2370

setting the given values, we get

z = (5500-11830)/2370 = -6330/2370 = -2.671

We know that any z value which is less than -1.96 and greater than +1.96 is considered as unsual

So, calculated z score is -2.671 is less than -1.96

So, it is an unsual z score